OpenStudy (anonymous):

please help Given that f(x) = 4x − 3 and g(x) = quantity of 2x minus 1, divided by 3, solve for g(f(2)) 1 3 9 15

OpenStudy (anonymous):

@Abmon98

OpenStudy (anonymous):

@bass53

OpenStudy (anonymous):

sry have one more kind of stretching the time

OpenStudy (anonymous):

would it be 1

OpenStudy (abmon98):

f(x)=4x-3 g(x)=2x-1/3 g(f(2)) 2(4x-3)-1/3 2(4(2)-3)-1/3=2(8-3)-1/3=2(5)-1/3=10-1/3=9/3=3

OpenStudy (anonymous):

so 3

OpenStudy (anonymous):

b

OpenStudy (abmon98):

yes